Your stock solution of tetracycline is prepared by dissolving of 0.375g of tetracycline in 25 ml ethanol and 25 ml distilled water.
How much stock do you have to add to prepare 1 L of medium with 15 ug/ml of tetracycline?
Answer #1

This problem can be solved by applying the data through the equation C1V1 = C2V2

C1 = Initial concentration or concentration of stock solution

C2 = Final concentration or concentration of working solution

V1 = Initial volume or volume of stock to be added

V2 = Final volume or volume of working solution

Here,

C1 = 0.375/25+25 gm/ml =0.375/50 gm/ml = 375/50 mg/ml = 7.5 mg/ml = 7500 ug/ml

C2 = 15 ug/ml

V2 = 1 L = 1000 ml

V1 = ?

As, C1V1 = C2V2

Hence, V1 = C2V2/C1

So, V1 = 15×1000/7500 = 2 ml

So, 2 ml of stock has to be added to prepare 1 L of medium with 15 ug/ml tetracycline

**** If you like the answer please hit the thumbs-up button. Thank you :-)